OpenStudy (anonymous):

HELPPPPPPPPPPPPPPPPPP Which of the following options results in a graph that shows exponential decay?

OpenStudy (anonymous):

f(x) = 0.6(2)x f(x) = 3(0.7)x f(x) = 0.4(1.6)x f(x) = 20(3)x

OpenStudy (anonymous):

@MathGuyThatWantsHelp @welshfella @pooja195

OpenStudy (anonymous):

ANYONE

OpenStudy (anonymous):

@MathGuyThatWantsHelp @pooja195 @welshfella

OpenStudy (anonymous):

@CBARREDO1

OpenStudy (anonymous):

i would have to say the third one since when you do the math on the other ones it is equal or greater than 1

OpenStudy (anonymous):

thx

OpenStudy (cbarredo1):

use this app it's really helpful, it'll make the graph for you https://www.desmos.com/calculator

OpenStudy (cbarredo1):

f(x) = 0.4(1.6)x

OpenStudy (irishboy123):

none of these really show that much as your notation is unclear. perhaps you could draw them using the Draw button